# Optimizing Probabilistic Conformal Prediction with Vectorized Non-Conformity Scores

### v1

@@ Appendix D Implementation Details
For the estimation of conditional distribution for CQR and CHR, we implemented two base models, including quantile regression neural network (QNN) and quantile regression forest (QRF). We used QNN as the base model for CHR and CQR in the synthetic data experiments and QRF for the real data experiments. For the conditional probabilistic models for PCP and PCP-VCR, we conduct nearest neighbor regression (NN) for $1$-dimensional synthetic data and mixture density network (MDN) for $2$-dimensional synthetic data. For the real data experiments, we use QRF as the base models for PCP and PCP-VCR, the same with CHR and CQR.
@@ Appendix E Experiments Results
In this section, we present the additional experimental results for synthetic data. We present the visualizations of the prediction set provided by different baseline methods on S-shape, Circles, Spirals, Unbalanced clusters, and mixture-Gaussian datasets in Figure 6 and 7. For mixture-Gaussian data, we generate $X\sim\textup{Uniform}(0,1)$, and let $Y|X\sim 0.7\textup{N}((X,0)^{T},I)+0.3\textup{N}((5+X,0)^{T},I)$ for identity matrix $I\in\mathbb{R}^{2\times 2}$. In the right subplot of Figure 7, we present the comparison of the prediction set of PCP and PCP-VCR. A key observation here is that in the low-density region (right component), PCP-VCR has a smaller radius for almost every circle, whereas in the high-density region (left component), PCP-VCR has a larger region covered. This demonstrates the ability of PCP-VCR to adapt to the local density by leveraging ranked samples.
Here, we also investigate the effect of $K$ on coverage efficiency. We provide an example in Figure 8, which presents the ratio of PCP prediction set length over PCP-VCR prediction set length for S-shape data under different choices of $K=3,5,10,15,20$ and different numbers of calibration data. The curve plot represents the mean, and the shaded area indicates the standard error of the ratio over 100 experiments. Under all settings, the ratio consistently remains above 1, showing that PCP-VCR provides a more efficient prediction set. Additionally, the ratio decreases as $K$ increases, suggesting that the benefit of PCP-VCR decreases with larger $K$.

### v2

@@ Appendix D Implementation Details
For the estimation of conditional distribution for CQR and CHR, we implemented two base models, including quantile regression neural network (QNN) and quantile regression forest (QRF). We used QNN as the base model for CHR and CQR in the synthetic data experiments and QRF for the real data experiments. For the conditional probabilistic models for PCP and GCP-VCR, we conduct nearest neighbor regression (NN) for $1$-dimensional synthetic data and mixture density network (MDN) for $2$-dimensional synthetic data. For the real data experiments, we use QRF as the base models for PCP and GCP-VCR, the same with CHR and CQR.
@@
Figure 7: Comparison of prediction set ($\alpha=0.1$) generated by different methods on $2$-dimensional synthetic mixture Gaussian data. The mixture Gaussian data is generated with 2 components with weights 0.7 and 0.3 from left to right, respectively. The blue shade indicates the true conditional density. In the left subplot, we present the rectangle-shaped prediction regions generated by CQR, CHR and R2CCP. In the right subplot, we present the circle-shaped prediction regions provided by GCP-VCR and PCP.
Figure 8: Ratio of PCP prediction set length over GCP-VCR prediction set length under different $K=3,5,10,15,20$ and number of calibration data for S-shape data. The blue curve represents the mean, and the shaded area indicates the standard error of the ratio. We observe that the ratio remains consistently above 1, demonstrating that GCP-VCR achieves better efficiency.
Figure 9: Comparison of miscoverage rate of PCP and GCP-VCR for MNIST data under different $K$. The blue and red bars represent PCP and GCP-VCR results, respectively, with black lines showing the standard error of the miscoverage rate across 30 experiments. GCP-VCR consistently achieves a lower miscoverage rate than PCP, indicating a more efficient prediction set.
@@ Appendix E Experiments Results
In this section, we present the additional experimental results for synthetic data and MNIST data.
E.1 Synthetic Data Results
We present the visualizations of the prediction set provided by different baseline methods on S-shape, Circles, Spirals, Unbalanced clusters, and mixture-Gaussian data in Figure 6 and 7. For mixture-Gaussian data, we generate $X\sim\textup{Uniform}(0,1)$, and set $Y|X\sim 0.7\textup{N}((X,0)^{T},I)+0.3\textup{N}((5+X,0)^{T},I)$ for identity matrix $I\in\mathbb{R}^{2\times 2}$. In the right subplot of Figure 7, we present the comparison of the prediction set of PCP and GCP-VCR. A key observation here is that in the low-density region (right component), GCP-VCR has a smaller radius for almost every circle, whereas in the high-density region (left component), GCP-VCR has a larger region covered. This demonstrates the ability of GCP-VCR to adapt to the local density by leveraging ranked samples.
Here, we also investigate the effect of $K$ on coverage efficiency. We provide an example in Figure 8, which presents the ratio of PCP prediction set length over GCP-VCR prediction set length for S-shape data under different choices of $K=3,5,10,15,20$ and different numbers of calibration data. The curve plot represents the mean, and the shaded area indicates the standard error of the ratio over 100 experiments. Under all settings, the ratio consistently remains above 1, showing that GCP-VCR provides a more efficient prediction set. Additionally, the ratio decreases as $K$ increases, suggesting that the benefit of GCP-VCR decreases with larger $K$.
E.2 MNIST Data Results
In Figure 9, we present the mean and standard error of the miscoverage rate of PCP and GCP-VCR for MNIST experiment with different $K$. The blue and red bars indicate the miscoverage rate of PCP and GCP-VCR, respectively. The dark line shows the standard error of miscoverage rate over 30 experiments. We observe that GCP-VCR consistently provides a lower miscoverage rate than PCP, indicating a higher efficiency of the prediction set. This also explains why the random images of GCP-VCR aligns more closely with the target as shown in Figure 3.